OpenStudy (anonymous):

Identify the vertex focus and directrix of the graph x^2-8x-28y-124=0 a. vertex (4,5) focus (4,2) directrix y=2. b.vertex(-4,5) focus (0,7) directrix y=7 c.vertex (4,-5) focus (4,2) directrix y=-12 d.vertex (-4,5) focus (4,-12) directrix y=2

OpenStudy (campbell_st):

this is \[x^2 - 8x + 16 = 28y + 140\] or \[(x -4)^2 = 28(y + 5)\] \[ (x^-4)^2 = 4 \times 7 (y +5)\] the equation is now in the form \[(x+h)^2 = 4a(y+k)\] the parabola is concave up h = 4, k = -5, a = 7 vertex is at (h, k) focus is (h, k + a) directrix is y = k - a